The north wall of a barn is a rectangle 30 feet high and 66 feet long. There are three windows in the wall, each 5 by 7 feet. If a gallon of paint will cover 375 ft2, how many gallons of paint must the painter buy to paint the wall?
Answer by Cromlix(4381)

Area of North Wall = 30 * 66 = 1980 ft^2
Area of window = 5 * 7 = 35 ft^2
3 windows = 105 ft^2
1980 - 105 = 1875 ft^2
A gallon of paint will cover 375 ft^2
1875/375 = 5
So, 5 gallons of paint are required.
Hope this helps.
:-)